Known Issues

Title Cat. Res. Description Post Upd.sort ascending
System Downtime 9/29/13 Outage Resolved

OSC systems will be offline on September 29th, 2013 for maintenance. Please visit osc.edu/n for more information.

6 months 4 weeks ago 6 months 3 weeks
8AM 9/11/13 - Brief network disruption to reboot a switch Network Resolved

At 8AM on September 11, 2013, we will be rebooting a network switch to replace a failed card in the switch. Network will be disrupted for 10 to 15 minutes while the work is done. Filesystem mounts... (Read more)

7 months 2 weeks ago 6 months 4 weeks
Brief disruption of GPFS on 8/28/2013 filesystem Resolved

On the morning August 28th, 2013 we will briefly disrupt the GPFS filesystem to reboot servers. This is necessary to upgrade the GPFS system. The in-place upgrade should only briefly interrupt... (Read more)

7 months 4 weeks ago 7 months 4 weeks
Brief disruption on 8/1/2013 at 8AM Network Resolved

At 8AM on the morning of 8/1/2013, we will be replacing some faulty hardware in our network infrastructure. Unfortunately, this work cannot be delayed until the next downtime, and the replacement... (Read more)

8 months 3 weeks ago 8 months 2 weeks
Poor network performance on some filesystems filesystem Resolved

We are experiencing some network performance issues on a cluster of servers involved with providing GPFS and some project filesystems. GPFS appears to be functioning acceptably, but proj01, proj02... (Read more)

9 months 3 days ago 9 months 3 days
Network card re-seat Network Resolved

At 8AM on Tuesday, July 9th 2013, we will be re-seating a network card in a switch at our operations center. It is possible that a brief (~10 minute) outage may occur. Jobs will pause for the... (Read more)

9 months 2 weeks ago 9 months 2 weeks
6/4/13 Scheduled Downtime Outage Resolved

HPC systems are currently offline for scheduled maintenance. See osc.edu/n for more information.

10 months 3 weeks ago 10 months 3 weeks
Brief interruption of services for some users filesystem Resolved

Today, May 14 2013, at 12:45PM we will be temporarily removing one of the home directory servers from service to address some reliability issues. Users with home... (Read more)

11 months 2 weeks ago 11 months 2 weeks
Backups of /nfs/gpfs Backups Resolved

Changes to files on /nfs/gpfs may not be backed up during the following evening's backup, as would normally be expected. The backup software is attempting to recreate a full backup... (Read more)

1 year 14 hours ago 11 months 3 weeks
Brief interruption of batch services on 4/17 Batch Resolved

On April 17th 2013, at roughly 2PM, we will be rebooting the batch server on the Oakley cluster. Running jobs will not be affected, but there will be a brief disruption in scheduling, as well as... (Read more)

1 year 6 days ago 1 year 6 days

Pages